# Enumeration of maps with tight boundaries and the Zhukovsky transformation

### v1

@@ C.2.2 Expressing the moments in terms of derivatives of $\mathbf{Z}$
Now, let us rewrite the equations (3) defining $R$ and $S$ in the more symmetric form
@@
$Z_{0}(r,s)=r-\sum_{i\geq 2}t_{i}\sum_{l}\binom{i-1}{\frac{i-l}{2},\frac{i-l}{2}-1,l}r^{\frac{i-l}{2}}s^{l}$ (191)
@@
$Z_{1}(r,s)=s-\sum_{i\geq 2}t_{i}\sum_{l}\binom{i-1}{\frac{i-l-1}{2},\frac{i-l-1}{2},l}r^{\frac{i-l-1}{2}}s^{l}\,.$ (192)
Then we obtain that
$\bar{M}_{+,h}=R^{-1/2}Q^{[0]}_{h}(r\partial_{r})Z_{0}(r,s)|_{r=R,s=S}+Q^{[1]}_{h}(r\partial_{r})Z_{1}(r,s)|_{r=R,s=S}\,,$ (193)
and
$(-1)^{h}\bar{M}_{h,-}=R^{-1/2}Q^{[0]}_{h}(r\partial_{r})Z_{0}(r,s)|_{r=R,s=S}-Q^{[1]}_{h}(r\partial_{r})Z_{1}(r,s)|_{r=R,s=S}\,.$ (194)
By Leibniz’ formula, this means that the (renormalized) moments are (possibly with some factors $R^{-1/2}$) linear combinations of the expressions
$R^{k}\frac{\partial^{k}Z_{\epsilon}}{\partial r^{k}}(R,S)\,,\qquad 1\leq k\leq h+1\,,\quad\epsilon\in\{0,1\}.$ (195)
@@ C.2.3 Structure of derivatives of $\mathbf{Z}$
To this end, we need to understand better the partial derivatives of $R,S$ with respect to $t_{0},t_{1}$. By applying a simple induction argument based on Proposition B.3 (and noting that the differential operators $D_{\ell}$ for $\ell\in\{0,1\}$ are given by $D_{\ell}=\partial/\partial t_{\ell}$), we obtain the following statement.
@@
One has
$\frac{\partial^{a+b}R}{\partial t_{0}^{a}\partial t_{1}^{b}}=(a+b)!R^{\frac{b}{2}+1}P_{0}^{(a,b)}\left(\frac{R^{(k)}}{R},\frac{S^{(k)}}{\sqrt{R}},1\leq k\leq a+b\right)$ (196)
and
$\frac{\partial^{a+b}S}{\partial t_{0}^{a}\partial t_{1}^{b}}=(a+b)!R^{\frac{b+1}{2}}P_{1}^{(a,b)}\left(\frac{R^{(k)}}{R},\frac{S^{(k)}}{\sqrt{R}},1\leq k\leq a+b\right)$ (197)
where $P^{(a,b)}_{0},P^{(a,b)}_{1}$ are polynomials of valuation [^9] In Eynard’s book, this would be coined as “homogeneous, where $R^{(k)}$ and $S^{(k)}$ are considered of degree $k$”. $a+b$, where a derivative of order $k$ counts for $k$ in the valuation. The normalization $(a+b)!$ is here to make things more convenient afterwards.
Note that these derivatives are of the form $T^{(0)}_{1,\ldots,1,0,\ldots 0}$ in both cases, with $a$ terms equal to 00 and $b+2$ terms equal to $1$ in the first case, and with $a+1$ terms equal to 00 and $b+1$ terms equal to $1$ in the second case, so in fact we have $P_{0}^{(a+1,b-1)}=P_{1}^{(a,b)}$.
@@
$\frac{\partial^{k}Z_{\epsilon}}{\partial r^{k}}(\mathbf{U}(\mathbf{t}))=\sum_{\tau\in\mathcal{T}^{\mathbf{0}}_{\epsilon}(k)}\prod_{v\in\tau}(d_{\mathbf{t}}\mathbf{U})^{-1}_{i_{v},j_{v}}\prod_{v\in\tau,k_{v}(\tau)>0}\frac{1}{k_{v}(\tau)!}\frac{\partial^{k_{v}(\tau)}U_{j_{v}}}{\partial t_{i_{v1}}\cdots\partial t_{i_{vk_{v}(\tau)}}}(\mathbf{t})$ (198)
@@
where $\mathcal{D}(i,j)$ is $R^{\prime}/R$ if $i+j$ is even, and $-S^{\prime}/\sqrt{R}$ otherwise. Hence, the contribution of edges of $\tau$—the first product in (198)—is
$\prod_{v\in V(\tau)}R^{-1-\frac{i_{v}}{2}+\frac{j_{v}}{2}}\frac{\mathcal{D}(i_{v},j_{v})}{\left(\frac{R^{\prime}}{R}\right)^{2}-\left(\frac{S^{\prime}}{\sqrt{R}}\right)^{2}}\,,$ (201)
@@
$\prod_{v\in\tau:k_{v}(\tau)>0}R^{\frac{\sum_{l=1}^{k_{v}(\tau)}i_{vl}}{2}+1-\frac{j_{v}}{2}}P_{j_{v}}^{\left(\sum_{l=1}^{k_{v}(\tau)}(1-i_{vl}),\sum_{l=1}^{k_{v}(\tau)}i_{vl}\right)}\left(\frac{R^{(l)}}{R},\frac{S^{(l)}}{\sqrt{R}},1\leq l\leq k_{v}(\tau)\right)\,.$ (202)
@@
$-k-i_{\varnothing}/2+\sum_{v\in\tau:k_{v}(\tau)=0}j_{v}/2=-k-\epsilon/2\,,$ (203)
since $\tau\in\mathcal{T}^{\mathbf{0}}_{\epsilon}(k)$. The rest of the contribution of $\tau$ is a polynomial of valuation $12\#\tau-$ 1 times $((R^{\prime}/R)^{2}-(S^{\prime}/\sqrt{R})^{2})^{-\#\tau}$. Multiplying and dividing by the determinant of $d_{\mathbf{t}}\mathbf{U}$ to the power $12k+$ 1, We finally obtain a contribution of $\tau$ of the form
$R^{-k-\epsilon/2}\frac{\Pi^{\tau}_{\pm,k}\left(\frac{R^{(l)}}{R},\frac{S^{(l)}}{\sqrt{R}},1\leq l\leq k\right)}{\left(\left(\frac{R^{\prime}}{R}\right)^{2}-\left(\frac{S^{\prime}}{\sqrt{R}}\right)^{2}\right)^{2k+1}}\,,$ (204)
where $\Pi^{\tau}_{\pm,k}$ are polynomials of valuation $14k+$ 1. After multiplying by $R^{k}$, summing over all possible $\tau\in\mathcal{T}^{\mathbf{0}}_{\epsilon}(k)$, and applying (193) and the remark below the latter formula, we finally obtain that
@@
[…]
@@ Appendix D Higher order differentials of inverse functions of several variables
Here we provide a general discussion revisiting Lagrange inversion that is used in Section C.2.2. The results are probably known in some other form, the paper by Warren Johnson cited in [budd2020irreducible] pointing as far as Sylvester.
@@
where $a_{1},-a_{2},-a_{3},\ldots$ are, up to a multiplicative coefficient, the successive differentials of $\mathbf{f}$ at 00, which are respectively symmetric $k$-linear, with the explicit expression $a_{1}(\mathbf{x})=\sum_{i=1}^{n}x_{i}\partial\mathbf{f}/\partial x_{i}(0)$, and
$-a_{k}(\mathbf{x},\ldots,\mathbf{x})=\frac{1}{k!}\sum_{i_{1},i_{2},\ldots,i_{k}\in\{1,\ldots,n\}}\frac{\partial^{k}\mathbf{f}}{\partial x_{i_{1}}\ldots\partial x_{i_{k}}}(0)x_{i_{1}}\cdots x_{i_{k}}\\ =\sum_{a_{1}+\cdots+a_{n}=k}\frac{\partial^{k}\mathbf{f}}{\partial x_{1}^{a_{1}}\cdots\partial x_{n}^{a_{n}}}(0)\frac{x_{1}^{a_{1}}}{a_{1}!}\cdots\frac{x_{n}^{a_{n}}}{a_{n}!}\,.$ (208)
@@
$g_{\epsilon}(\mathbf{y})=\sum_{\tau\in\mathcal{T}_{\epsilon}}\prod_{v\in\tau}(a_{1}^{-1})_{i_{v},j_{v}}(a_{k_{v}(\tau)})_{j_{v}}^{i_{v1},i_{v2},\ldots,i_{vk_{v}(\tau)}}\,,$ (210)
with the convention that $(a_{0})_{j}^{\varnothing}=y_{j}$, that is, the contributions of the variable $\mathbf{y}$ comes from the leaves of the trees. The partial derivatives of order $k$ of $g_{\epsilon}$ come from the finite family $\mathcal{T}_{\epsilon}(k)$ of elements of $\mathcal{T}_{\epsilon}$ with $k$ leaves. More precisely, for a given composition $\mathbf{a}=(a_{1},\ldots,a_{n})$ of $k$ with $n$ parts, one has
$\frac{\partial^{k}g_{\epsilon}}{\partial x_{1}^{a_{1}}\cdots\partial x_{n}^{a_{n}}}(0)=\sum_{\tau\in\mathcal{T}^{\mathbf{a}}_{\epsilon}}\prod_{v\in\tau}(a_{1}^{-1})_{i_{v},j_{v}}\prod_{v\in\tau:k_{v}(\tau)>0}(a_{k_{v}(\tau)})_{j_{v}}^{i_{v1},i_{v2},\ldots,i_{vk_{v}(\tau)}}\,,$ (211)
where $\mathcal{T}^{\mathbf{a}}_{\epsilon}$ is the set of trees $\tau\in\mathcal{T}_{\epsilon}(k)$ with $a_{i}$ leaves $v$ labeled by $j_{v}=i$, for $1\leq i\leq n$.
@@
$\frac{\partial^{k}g_{\epsilon}}{\partial x_{1}^{a_{1}}\cdots\partial x_{n}^{a_{n}}}(\mathbf{f}(\mathbf{x}))=\sum_{\tau\in\mathcal{T}^{\mathbf{a}}_{\epsilon}(k)}\prod_{v\in\tau}(d_{\mathbf{x}}\mathbf{f}^{-1})_{i_{v},j_{v}}\prod_{v\in\tau:k_{v}(\tau)>0}\frac{-1}{k_{v}(\tau)!}\frac{\partial^{k_{v}(\tau)}f_{j_{v}}}{\partial x_{i_{v1}}\cdots\partial x_{i_{vk_{v}(\tau)}}}(\mathbf{x})\,.$ (212)
[…]

### v2

@@ C.2.2 Relating the moments with the recursions determining $R,S$
The expression (184) for the moments can be compared to the equations (3) defining $R$ and $S$, which we may write in the more compact and symmetric form
@@
$Z_{0}(r,s)=r-\sum_{i\geq 2}t_{i}\sum_{l=i\,\mathrm{mod}\,2}\binom{i-1}{\frac{i-l}{2},\frac{i-l}{2}-1,l}r^{\frac{i-l}{2}}s^{l}$ (194)
@@
$Z_{1}(r,s)=s-\sum_{i\geq 2}t_{i}\sum_{l=i+1\,\mathrm{mod}\,2}\binom{i-1}{\frac{i-l-1}{2},\frac{i-l-1}{2},l}r^{\frac{i-l-1}{2}}s^{l}\,.$ (195)
Combining this with (184), we obtain that, for every $h\geq 1$,
$(\pm 1)^{h}\bar{M}_{\pm,h}=-\sum_{\epsilon\in\{0,1\}}(\pm 1)^{\epsilon}R^{\frac{\epsilon-1}{2}}Q_{h}^{[\epsilon]}(r\partial_{r})Z_{\epsilon}(r,s)|_{r=R,s=S}\,$, (196)
where we note that the term $r$ in (194) is annihilated by the differential operator $Q_{h}^{[0]}(r\partial_{r})$ thanks to (185). By Leibniz’s formula, this means that the renormalized moments are linear combinations of the expressions
$R^{k+\frac{\epsilon-1}{2}}\frac{\partial^{k}Z_{\epsilon}}{\partial r^{k}}(R,S)\,,\qquad 1\leq k\leq h+1\,,\quad\epsilon\in\{0,1\}.$ (197)
@@ C.2.3 Structure of derivatives of $\mathbf{Z}$
To this end, we need to better understand the partial derivatives of $R,S$ with respect to $t_{0},t_{1}$. We recall that $R^{(k)}$ and $S^{(k)}$ correspond to the derivatives with respect to $t=t_{0}$.
@@
There exists a family of polynomials $P^{(a,b)}(x_{k},y_{k},1\leq k\leq a+b)$ such that, for any integers $a,b\geq 0$, we have
$\frac{\partial^{a+b}R}{\partial t_{0}^{a}\partial t_{1}^{b}}=R^{\frac{b}{2}+1}P^{(a,b)}\left(\frac{R^{(k)}}{R},\frac{S^{(k)}}{\sqrt{R}},1\leq k\leq a+b\right)$ (198)
and, if moreover $(a,b)\neq(0,0)$,
$\frac{\partial^{a+b}S}{\partial t_{0}^{a}\partial t_{1}^{b}}=R^{\frac{b+1}{2}}P^{(a+1,b-1)}\left(\frac{R^{(k)}}{R},\frac{S^{(k)}}{\sqrt{R}},1\leq k\leq a+b\right)\,.$ (199)
It is defined by induction as follows: we have $P^{(a,0)}=x_{a}$ for every $a\geq 0$, with the convention that $x_{0}=1$, $P^{(a+1,-1)}=y_{a}$ for every $a\geq 1$, and, for every $a,b\geq 0$,
$P^{(a,b+1)}=\left(\left(\frac{b}{2}+1\right)y_{1}+\sum_{k=1}^{a+b}\left(\sum_{i=0}^{k-1}\binom{k}{i}x_{i}y_{k+1-i}\partial_{x_{k}}+\left(x_{k+1}-\frac{y_{1}y_{k}}{2}\right)\partial_{y_{k}}\right)\right)P^{(a,b)}\,.$ (200)
Moreover, the polynomial $P^{(a,b)}(\bar{x}_{k}^{k},\bar{y}_{k}^{k},1\leq k\leq a+b)$ in the variables $\bar{x}_{k},\bar{y}_{k}$ is homogeneous of degree $a+b$.
Proof.
We first make the trivial observation that for $b=0$, (198) holds with $P^{(a,0)}=x_{a}$ for $a\geq 0$, and (199) holds with $P^{(a+1,-1)}=y_{a}$ for $a\geq 1$. Moreover, by the combinatorial definition of $R$ and $S$ given in Section 1.2, the derivatives in (198) and (199) are of the form $T^{(0)}_{1,\ldots,1,0,\ldots 0}$ in both cases, with $a$ terms equal to 0 and $b+2$ terms equal to $1$ in the first case, and with $a+1$ terms equal to 0 and $b+1$ terms equal to $1$ in the second case. This shows that, for $a\geq 0$ and $b\geq 1$, (199) is nothing but (198) applied to $a+1$ and b1b-1 instead of $a,b$. Furthermore, we deduce from this discussion and Theorem A.1 the identities
$\frac{\partial R}{\partial t_{1}}=RS^{\prime}\,,\qquad\frac{\partial S}{\partial t_{1}}=R^{\prime}\,$, (201)
which are consistent with the particular values $P^{(0,1)}=y_{1}$ and $P^{(1,0)}=x_{1}$.
In view of these preliminary remarks, it now suffices to establish (198) for general $b$, which we will do by induction. Assume that it holds at rank $b\geq 0$, and let us establish it at rank $b+1$. To this end, we write
$\frac{\partial^{a+b+1}R}{\partial t_{0}^{a}\partial t_{1}^{b+1}}=\frac{\partial}{\partial t_{1}}\left(R^{\frac{b}{2}+1}P^{(a,b)}\left(\frac{R^{(k)}}{R},\frac{S^{(k)}}{\sqrt{R}},1\leq k\leq a+b\right)\right)\,$, (202)
and we expand the right-hand side using the Leibniz and chain rules. The first term in the expansion, obtained by differentiating $R^{b/2+1}$, is equal to
$\left(\frac{b}{2}+1\right)R^{\frac{b}{2}}\frac{\partial R}{\partial t_{1}}P^{(a,b)}=\left(\frac{b}{2}+1\right)R^{\frac{b+1}{2}+1}\frac{S^{\prime}}{\sqrt{R}}P^{(a,b)}\,.$ (203)
The term obtained by differentiating with respect to the variable $x_{k}=R^{(k)}/R$ in $P^{(a,b)}$ is equal to
$R^{\frac{b}{2}+1}\partial_{x_{k}}P^{(a,b)}\frac{\partial}{\partial t_{1}}\left(\frac{R^{(k)}}{R}\right)\,$, (204)
where we may write
$\begin{split}\frac{\partial}{\partial t_{1}}\left(\frac{R^{(k)}}{R}\right)&=\frac{\frac{\partial R^{(k)}}{\partial t_{1}}}{R}-\frac{\frac{\partial R}{\partial t_{1}}R^{(k)}}{R^{2}}=\frac{\frac{\partial^{k}(RS^{\prime})}{\partial t_{0}^{k}}}{R}-\frac{S^{\prime}R^{(k)}}{R}\\ &=\sqrt{R}\sum_{i=0}^{k-1}\binom{k}{i}\frac{R^{(i)}}{R}\frac{S^{(k+1-i)}}{\sqrt{R}}\end{split}$ (205)
by (201) and Leibniz’s formula. Finally, the term obtained by differentiating with respect to the variable $y_{k}=S^{(k)}/\sqrt{R}$ in $P^{(a,b)}$ is equal to
$R^{\frac{b}{2}+1}\partial_{y_{k}}P^{(a,b)}\frac{\partial}{\partial t_{1}}\left(\frac{S^{(k)}}{\sqrt{R}}\right)\,$, (206)
where we may write
$\frac{\partial}{\partial t_{1}}\left(\frac{S^{(k)}}{\sqrt{R}}\right)=\frac{\frac{\partial S^{(k)}}{\partial t_{1}}}{\sqrt{R}}-\frac{\frac{\partial R}{\partial t_{1}}S^{(k)}}{2R^{3/2}}=\sqrt{R}\left(\frac{R^{(k+1)}}{R}-\frac{1}{2}\frac{S^{\prime}}{\sqrt{R}}\frac{S^{(k)}}{\sqrt{R}}\right)\,.$ (207)
Putting (203)–(207) together, we obtain that (198) is valid for $(a,b+1)$, where $P^{(a,b+1)}$ is given by (200), as wanted. The statement about the degree and homogeneity of $P^{(a,b)}(\bar{x}_{k}^{k},\bar{y}_{k}^{k},1\leq k\leq a+b)$ is readily derived by induction, using (200). ∎
@@
$\frac{\partial^{k}Z_{\epsilon}}{\partial r^{k}}(\mathbf{U}(\mathbf{t}))=\sum_{\tau\in\mathcal{T}^{\mathbf{0}}_{\epsilon}(k)}\prod_{v\in\tau}(d_{\mathbf{t}}\mathbf{U})^{-1}_{i_{v},j_{v}}\prod_{v\in\tau:k_{v}(\tau)>0}\frac{1}{k_{v}(\tau)!}\frac{\partial^{k_{v}(\tau)}U_{j_{v}}}{\partial t_{i_{v1}}\cdots\partial t_{i_{vk_{v}(\tau)}}}(\mathbf{t})$ (208)
@@
where $\mathcal{D}(i,j)$ is $R^{\prime}/R$ if $i+j$ is even, and $S^{\prime}/\sqrt{R}$ otherwise. Hence, the contribution of edges of $\tau$—the first product in (208)—is
$\prod_{v\in\tau}R^{-1-\frac{i_{v}}{2}+\frac{j_{v}}{2}}\frac{\mathcal{D}(i_{v},j_{v})}{\left(\frac{R^{\prime}}{R}\right)^{2}-\left(\frac{S^{\prime}}{\sqrt{R}}\right)^{2}}\,$, (211)
@@
$\prod_{v\in\tau:k_{v}(\tau)>0}\frac{R^{\frac{\sum_{l=1}^{k_{v}(\tau)}i_{vl}}{2}+1-\frac{j_{v}}{2}}}{k_{v}(\tau)!}P^{\left(\sum_{l=1}^{k_{v}(\tau)}(1-i_{vl})+j_{v},\sum_{l=1}^{k_{v}(\tau)}i_{vl}-j_{v}\right)}\left(\frac{R^{(l)}}{R},\frac{S^{(l)}}{\sqrt{R}},1\leq l\leq k_{v}(\tau)\right)\,.$ (212)
@@
$k-i_{\varnothing}/2+\sum_{v\in\tau:k_{v}(\tau)=0}j_{v}/2=-k-\epsilon/2\,$, (213)
since $\tau\in\mathcal{T}^{\mathbf{0}}_{\epsilon}(k)$. The rest of the contribution of $\tau$ is a homogeneous polynomial of degree $2\#\tau-1$ (in the sense of Proposition C.4, that is, where the variables of index $l$ are considered to have degree $l$), times $((R^{\prime}/R)^{2}-(S^{\prime}/\sqrt{R})^{2})^{-\#\tau}$. Note that $\tau$ is maximal and equal to 2k12k-1 when $\tau$ is a binary tree. Hence, we reduce the sum in the right-hand side of (208) to a common denominator by multiplying and dividing the contribution of $\tau$ by $((R^{\prime}/R)^{2}-(S^{\prime}/\sqrt{R})^{2})^{2k-1-\#\tau}$, which is thus of the form
$R^{-k-\epsilon/2}\frac{\Pi^{\tau}_{\pm,k}\left(\frac{R^{(l)}}{R},\frac{S^{(l)}}{\sqrt{R}},1\leq l\leq k\right)}{\left(\left(\frac{R^{\prime}}{R}\right)^{2}-\left(\frac{S^{\prime}}{\sqrt{R}}\right)^{2}\right)^{2k-1}}\,$, (214)
where $\Pi^{\tau}_{\pm,k}$ is a homogeneous polynomial of degree 4k34k-3. After multiplying by $R^{k}$, summing over all possible $\tau\in\mathcal{T}^{\mathbf{0}}_{\epsilon}(k)$, and applying (196) and the remark leading to (197), we finally obtain that
@@
where $\Pi_{\pm,h}$ are homogeneous polynomials of degree $4h+1$. Coming back to the “dimensionless” quantities $R^{h/2}M_{\pm,h}=\bar{M}_{\pm,h}/(\sqrt{R}M_{\pm,0})$, those are, recalling (176),
@@ Appendix D Higher order differentials of inverse functions of several variables
Here we provide a general discussion revisiting Lagrange inversion that is used in Section C.2.3. The results are probably known in some other form, the paper by Warren Johnson cited in [Bud22] pointing as far as Sylvester.
@@
where $a_{1},-a_{2},-a_{3},\ldots$ are, up to a multiplicative coefficient, the successive differentials of $\mathbf{f}$ at $\mathbf{0}$, which are respectively $k$-linear symmetric. Using tensor notations, we explicitly have
$(a_{1})_{i}^{j}=\frac{\partial f_{i}}{\partial x_{j}}(\mathbf{0})\,,\qquad(a_{k})_{i}^{j_{1},\ldots,j_{k}}=-\frac{1}{k!}\frac{\partial^{k}f_{i}}{\partial x_{j_{1}}\ldots\partial x_{j_{k}}}(\mathbf{0})\,,\quad k\geq 2\,.$ (218)
@@
$g_{\epsilon}(\mathbf{y})=\sum_{\tau\in\mathcal{T}_{\epsilon}}\prod_{v\in\tau}(a_{1}^{-1})_{i_{v}}^{j_{v}}(a_{k_{v}(\tau)})_{j_{v}}^{i_{v1},i_{v2},\ldots,i_{vk_{v}(\tau)}}\,$, (220)
with the convention that $(a_{0})_{j}^{\varnothing}=y_{j}$, that is, the contributions of the variable $\mathbf{y}$ come from the leaves of the trees. The partial derivatives of order $k$ of $g_{\epsilon}$ come from the finite family $\mathcal{T}_{\epsilon}(k)$ of elements of $\mathcal{T}_{\epsilon}$ with $k$ leaves. More precisely, for a given composition $\mathbf{c}=(c_{1},\ldots,c_{n})$ of $k$ with $n$ parts, one has
$\frac{\partial^{k}g_{\epsilon}}{\partial x_{1}^{c_{1}}\cdots\partial x_{n}^{c_{n}}}(\mathbf{0})=\sum_{\tau\in\mathcal{T}^{\mathbf{c}}_{\epsilon}}\prod_{v\in\tau}(a_{1}^{-1})_{i_{v}}^{j_{v}}\prod_{v\in\tau:k_{v}(\tau)>0}(a_{k_{v}(\tau)})_{j_{v}}^{i_{v1},i_{v2},\ldots,i_{vk_{v}(\tau)}}\,$, (221)
where $\mathcal{T}^{\mathbf{c}}_{\epsilon}$ is the set of trees $\tau\in\mathcal{T}_{\epsilon}(k)$ with $c_{i}$ leaves $v$ labeled by $j_{v}=i$, for $1\leq i\leq n$.
@@
$\frac{\partial^{k}g_{\epsilon}}{\partial x_{1}^{c_{1}}\cdots\partial x_{n}^{c_{n}}}(\mathbf{f}(\mathbf{x}))=\sum_{\tau\in\mathcal{T}^{\mathbf{c}}_{\epsilon}(k)}\prod_{v\in\tau}(d_{\mathbf{x}}\mathbf{f}^{-1})_{i_{v}}^{j_{v}}\prod_{v\in\tau:k_{v}(\tau)>0}\frac{-1}{k_{v}(\tau)!}\frac{\partial^{k_{v}(\tau)}f_{j_{v}}}{\partial x_{i_{v1}}\cdots\partial x_{i_{vk_{v}(\tau)}}}(\mathbf{x})\,.$ (222)